OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 12.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

But isn't it solved a different way if it says "varies inversely" ?

OpenStudy (anonymous):

sorry, wait.

OpenStudy (anonymous):

y = k/3x 2 = k/3(6) k = 36 12 = 36/3x x = 1

OpenStudy (anonymous):

Thank you (: